the standard enthalpy of formation of liquid ethanol C2H5OH(I), is -227.63 kJ/mol. Calculate the heat of combustion at constant pressure and at constant volume of C2H5OH(I). ΔH°f, CO2= -393.5 kJ/mol; ΔH°f, H2O= -285.4 kJ/mol

C2H5OH(I)+3 O2(g) = 2 CO2(g) + 3 H2O


1
Expert's answer
2022-01-13T08:20:03-0500

ΔH°f, C2H5OH(I) = −227.63 kJ/mol

ΔH°f, O2(g) = 0 (simple substance)

ΔH°f, CO2(g) = −393.5 kJ/mol

ΔH°f, H2O(l) = −285.4 kJ/mol


Solution:

Balanced chemical equation:

C2H5OH(I) + 3O2(g) → 2CO2(g) + 3H2O(l)

To find the heat (ΔH°r) of the ethanol combustion reaction, use the formula for the standard enthalpy change of formation:

ΔH°r = ∑ΔH°f (products) − ∑ΔH°f (reactants)

ΔH°r = 3 × ΔH°f, H2O(l) + 2 × ΔH°f, CO2(g) − 3 × ΔH°f, O2(g) − ΔH°f, C2H5OH(I)

ΔH°r = 3 × (−285.4 kJ/mol) + 2 × (−393.5 kJ/mol) − 3 × 0 − (−227.63 kJ/mol) = −1415.57 kJ/mol

ΔH°r = −1415.57 kJ/mol


Answer: −1415.57 kJ/mol